# Steering LLMs for Formal Theorem Proving

### v4

@@ 1 Introduction
    Activation steering framework yielding consistent gains in proof generation quality on MiniF2F and PutnamBench.
This study not only improves how models reason about specific tasks, but also provides a practical, interpretable tool for steering reasoning. These learning’s provide crucial insights to interpret and understand LLMs for theorem proving better.
2 Related Work
Automatic Theorem Proving and Auto-formalization Prior work on automatic theorem proving typically follows the GPT-f framework (Polu and Sutskever,, 2020), which trains a language model on proof state and next tactic pairs, then proves theorems using this model with best-first tree search. Subsequent work has explored several directions, including data augmentation (Han et al.,, 2022), synthetic data for training language models (Rotella et al.,, 2025; Wang and Deng,, 2020), novel proof search methods (Wang et al.,, 2023), further training through curriculum learning (Polu et al.,, 2022), retrieval augmented generation (Yang et al.,, 2023), and practical tools (Welleck and Saha,, 2023).
Figure 1: In this example we show how steering vectors introduce natural language grounding in the proof improving the quality. We have highlighted the effect of steering vectors in yellow. The lines highlighted cyan represent incorrect tactics leading to incorrect proof.
A parallel thread focusing on improving automated theorem proving has explored auto-formalization (Wu et al.,, 2022), which incorporates informal mathematics into formal theorem proving. This direction recognizes the significant value of connecting informal mathematical intuition with formal proof systems. Draft-Sketch-Prove (Jiang et al.,, 2023) demonstrates that language models can leverage informal proofs to enhance formal proving abilities through few-shot prompting rather than supervised fine-tuning. Similarly, LeanStar (Lin et al.,, 2024) introduces ”thought-augmented reasoning” in each proof step, using informal reasoning to guide tactic prediction. However, approaches like LeanStar typically require supervised fine-tuning of large-language models on synthetic data to improve reasoning capabilities.
Our work addresses the fundamental challenge of understanding the relationship between informal and formal mathematics by learning task of producing thoughts directly in activation space to improve theorem proving at inference time. This approach provides a lightweight alternative to learning informal reasoning patterns through fine-tuning and offers insights into how language models internally represent and connect informal mathematical reasoning with formal proof steps. Understanding this relationship is crucial for developing more effective theorem proving systems that can leverage the complementary strengths of both informal mathematical intuition and formal verification.
Language model representation Our approach is inspired by previous research that shows that it is possible to obtain vector representation of tasks that language model learns in context (Hendel et al.,, 2023; Todd et al.,, 2024). These studies are part of broader body of work that examines the linear representation of features such as truthfulness (Azaria and Mitchell,, 2023; Li et al.,, 2023; Marks and Tegmark,, 2024), sentiment (Tigges et al.,, 2024), harmlessness (Zou et al.,, 2023; Zheng et al.,, 2024), sychophancy (Perez et al.,, 2023; Panickssery et al., 2024a, ; Sharma et al.,, 2024), factual knowledge (Gurnee and Tegmark,, 2024), and refusal (Arditi et al.,, 2024). In addition, recent works have employed sparse autoencoders to identify feature directions in an unsupervised manner (Bricken et al.,, 2023; Huben et al.,, 2024; Templeton et al.,, 2024).
A shared hypothesis across these works is that LLMs represent features or concepts as linear directions in activation space (Mikolov et al.,, 2013; Elhage et al.,, 2021; Nanda et al.,, 2023; Park et al.,, 2024). While recent studies suggest that not all features may be linearly encoded (Engels et al.,, 2025), the linearity assumption has been effective for both concept erasure and model steering. In our work we explore how the concept of generating natural language informal data is a linearly representable concept in activation space which improves the performance of these models to prove formal theorems.
Mechanistic Interpretability and Activation Patching Previous work has focused on localizing and editing factual associations within transformers (Meng et al.,, 2022) and probing hidden representations for high-level knowledge (Li et al.,, 2024; Dong et al.,, 2023). Such studies show implicit evaluations the ability of the model, complementing explicit benchmarks (Dong et al.,, 2023). A key technique in mechanistic interpretability is activation patching (Vig et al.,, 2020; Variengien and Winsor,, 2023), which modifies model activations in the residual stream at the time of inference to influence outputs (Zhao et al.,, 2025). Residual streams is the high-dimensional vector that carries and accumulates information across layers via skip (residual) connections allowing each layer to iteratively refine and add shared representation used for prediction. In figure 2 we demonstrate the process of generating activation vectors on the proof step generation task to enable language models to generate natural language patch that contains context for the next tactic.
For theorem proving, mechanistic interpretability should provide insight onto how well language models can comprehend natural language information and use it to comprehend formal proofs. By dissecting these representations, we can identify failure cases, refine step generation, and enhance the proof generation capability. We observe that effective steering transforms activations to align the model’s responses to a structured format to generate natural language information while writing a proof step that allows them to be better at proofing theorems.
3 Methodology
3.1 Model Choice
We build steering datasets for 7B parameter Llemma (Azerbayev et al.,, 2024), InternLM2 (Ying et al.,, 2024) and InternLM2.5-StepProver (Wu et al.,, 2024). These models are trained to generate formal theorems and on popular programming languages to improve their capability of reason about theorem’s better. InternLM family of models are based on LlaMA architecture (Touvron et al.,, 2023), however Llemma model is trained on CodeLlaMA architecture.
Figure 2: Steering Vectors are computed as difference of means for for $p^{+}$ and $p$
3.2 Tasks and Source Data
We constructed steering pairs from a randomly sampled subset $S$ of the Lean-STaR dataset  [^1] https://huggingface.co/datasets/ScalableMath/Lean-STaR-plus (Lin et al.,, 2024). This subset consists of $10,000$ unique data points. Each of these data points consist of a formal proof step and an informal thought required to generate that step.
Every natural prompt $p\in S$ is a formal proof step of a proof. We then create $p^{+}$ by inserting explicit reasoning step of natural language explanations of the proof state. This process gives us prompts pairs that are required to generate steering vectors.
After generating prompt pairs $(p,p^{+})$, we prompt the model to with both the prompts and collect responses $r_{p}$ and $r_{p^{+}}$. We then filter a subset $s\subset S$ where both $r_{p}$ and $r_{p^{+}}$ are valid proof steps and $r_{p}\neq r_{p^{+}}$. The resulting quadruples $\{p,p^{+},r_{p},r_{p^{+}}\}$ form the basis for steering vector construction.
Formal Theorem to Formal Proof Writing In this task, the model has generate a proof step-by-step for a given formal problem. We use standard benchmark for this task miniF2F (Zheng et al.,, 2022). This benchmark contains 244 test set theorems in Lean. To generate proof for the problem we use best first search as well as beam search. We describe the experiment setup in Section 4.
Informal Problem to Formal Proof Writing This task covers writing proof from informal problem statement. We use PUTNAMBENCH (Tsoukalas et al.,, 2024) as the benchmark for this task. This benchmark includes 657 problems in Lean and 640 problems in Coq. We use both these datasets to evaluate steering vectors.
3.3 Constructing Steering Vectors
To extract robust steering vectors, we first identify the activations where these vectors are linearly represented within the model. We focus on residual stream activations i.e the outputs of each transformer layer. Given an informal thought step we identify the layers where the activation vector is active at time of processing the last token.
To identify the causally relevant layers, we first extract steering vector candidate from every layer using Difference of Means method. This is a widely used technique (Turner et al.,, 2024; Arditi et al.,, 2024). This technique is based on constructing contrastive dataset that differs in a specific concept. For our task the contrastive dataset is $s$ and the concept we want to differentiate is producing NL information with the formal proof-step. We begin with a data set of pairs of prompts and their responses $\{(p_{i},p_{i}^{+},r_{p_{i}},r_{p_{i}^{+}})\}_{i=1}^{N}$, where $p_{i}^{+}$ includes informal reasoning steps and $p_{i}$ is the standard prompt. The model $M$ and extract its residual stream activation at the final token of layer $\ell$, denoted
$v_{p_{i}^{-}}^{\ell}=\mathrm{resid}(M(p_{i}^{-}),\ell)\quad\text{and}\quad v_{p_{i}^{+}}^{\ell}=\mathrm{resid}(M(p_{i}^{+}),\ell).$
To isolate the component capturing natural language information about the proof, we apply the Difference of Means method:
$u^{(\ell)}=\frac{1}{|s_{+}|}\sum_{p^{+}\in S_{+}}v_{p^{+}}^{\ell}\;-\;\frac{1}{|s_{-}|}\sum_{p^{-}\in S_{-}}v_{p^{-}}^{\ell},$
Where $v_{p}^{\ell}$ represents the activation vector of layer $\ell$ for a prompt $p$ and $s_{+}$ represents all the samples in the set $s$ that exhibit the targeted behavior while $s_{-}$ represent all the samples. In this scenario the Difference of Means vector is computed by subtracting overall mean activation from the mean activation of the behavior associated with the example. This isolates the direction in the activation space most associated with the target behavior while reducing the influence of general model biases.
Figure 3: Cosine similarity in InternLM2’s residual stream activations drops in later layers for informal reasoning queries, but gradually recovers.
4 Results and Analysis
Setup We construct proofs using best-first search method which is one of the most popular methods to evaluate the theorem proving ability of a language model (Polu and Sutskever,, 2020; Yang et al.,, 2023; Azerbayev et al.,, 2023; Lin et al.,, 2024). For a given language model $M$, we keep all unexpanded states $s_{i}$; each time we expand the best state $s_{i}$ and use the language model to sample $S$ net tactics $a_{i,1...S}$ for the current state $s_{i}$. Following standard practice (Polu and Sutskever,, 2020; Yang et al.,, 2023; Welleck and Saha,, 2023; Lin et al.,, 2024), we assume that the state with maximum negative log-probabilities is the ”best” state. Specifically, we select state $s_{i}$ with maximum $\sum_{j=0}^{i-1}-\log p(a_{j},s_{j})$, where $(s_{0},a_{0}),...,(s_{i-1},a_{i-1})$ is the proof trajectory before state $s_{i}$ and $\log p(a_{j},s_{j})$ is the average logarithmic probability of each generated token. We expand up to $N$ states and we get a successful proofs search when we reach any proof state with no goals.
We observed that the best first search perform poorly in our case as with the natural language tokens in our response the mean of log probabilities is noisy and inconsistent, making it harder to choose the post promising branch. This observation is consistent to Lean-STaR (Lin et al.,, 2024).
To overcome these limitation we apply Sampling technique proposed by Lean-STaR (Moura and Ullrich,, 2021). Sampling is based on sampling trajectories in parallel for each problem, each time generating at most N tactics. The illegal sampled tactics are ignored during sampling. This means at state $s_{i}$ if the sampled tactic is $a_{i}$ is only added to trajectory $(s_{0},a_{0}),(s_{1},a_{1}),...,(s_{i-1},a_{i-1})$ if $a_{i}$ is a legal tactic otherwise we sample another tactic from the model M.
This technique takes approximately same time on GPU as best-first search but accommodates natural language thoughts.
Figure 4: Effect of patched layers on proof generation in miniF2F-test using InternLM2. Y-axis: verified proofs; X-axis: number of patched layers.
Model Decoding N K S MiniF2F
LLEMMA-7B SEARCH 50 1 32 26.2%
InternLM2-7B SEARCH 50 1 32 30.3%
InternLM2-7B (SFT) SEARCH 50 1 32 30.7%
LEAN-COT (InternLM2-7B) SAMPLING 50 32 1 27.0%
LEAN-COT (InternLM2-7B) SEARCH 50 1 32 25.4%
LEAN-STAR (InternLM2-7B) SAMPLING 50 32 1 29.1%
LEAN-STAR (InternLM2-7B) SEARCH 50 1 32 26.2%
COPRA (WITH GPT-4) CUSTOMIZED - 60 1 29.9%
InternLM-2-Math-Plus SEARCH 100 1 32 43.4%
InternLM-2-StepProver SEARCH 100 1 32 48.8%
InternLM-2.5-StepProver SEARCH 600 2 32 50.7%
DeepSeekProver-V1.5 SEARCH 400 32 16 63.5%
InternLM-2.5-StepProver SEARCH 600 256 32 65.9%
LLEMMA-7B + Activation vector SAMPLING 50 32 1 28.1%
LLEMMA-7B + Activation vector SEARCH 50 1 32 26.3%
InternLM2-7B + Activation vector SAMPLING 50 32 1 32.4%
InternLM2-7B + Activation vector SEARCH 50 1 32 26.8%
InternLM-2.5-StepProver + Activation Vector SEARCH 600 2 32 49.18%
InternLM-2.5-StepProver + Activation Vector SAMPLING 600 2 32 52.4%
InternLM-2.5-StepProver + Activation Vector SEARCH 600 256 32 65.57%
InternLM-2.5-StepProver + Activation Vector SAMPLING 600 256 32 66.39%
Table 1: Pass rates on the miniF2F-test dataset with Lean. S is the number of tactics attempted at each expanded node (assumed to be 1 in sampling), and K is the total number of search or sampling attempts per problem.
Model Sampling Pass at
Problems Solved
Lean4 (Out of 657)
InternLM2 7B SEARCH 4096 4
InternLM2.5-StepProver SEARCH 2x32x600 6
Goedel-Prover-SFT SEARCH 512 7
ABEL SEARCH 596 7
Self-play Theorem Prover SEARCH 3200 8
Kimina-Prover-7B-Distill SEARCH 192 10
InternLM2 7B + Activation Steering SAMPLING 4096 4
InternLM2.5-StepProver + Activation Steering SAMPLING 2x32x600 7
Coq (out of 412)
InternLM2.5-StepProver SAMPLING 2x32x600 0
InternLM2.5-StepProver + Activation Steering SAMPLING 2x32x600 1
Table 2: Results on PutnamBench with Lean and Coq.
4.1 Main results
We construct distinct steering vectors for each model variant, enabling us to isolate and analyze the activation-level effects of distilled informal reasoning. A detailed comparison of activation patterns across models is presented in Appendix.
miniF2F Our results on the miniF2F test set are summarized in Table 1. Steering the model’s activations consistently improves performance across decoding strategies when compared to the base model. Notably, we observe substantial gains in pass rates under both best-first search and sampling strategies. One exception is a slight drop in performance for InternLM-2.5-StepProver under best-first search with activation steering. We hypothesize that this is due to the injected natural language content lowering the mean log-probability of tokens, which can affect deterministic decoding strategies more than stochastic ones.
PutnamBench Table 2 reports results on the Lean and Coq subsets of PutnamBench. Our findings further support the effectiveness of activation steering: despite the steering vectors being derived from Lean-based datasets, the augmented model successfully proves one Coq problem—an improvement over the base model’s zero success rate. This cross-system transfer underscores the generality and robustness of the distilled reasoning signal encoded in the steering vectors.
4.2 Research Questions
4.2.1 How effective and efficient are activation steering vectors?
Finetuning To understand the quality of our steering vectors, we do Low Rank Adapter fine-tuning (LORA) on InternLM2 on all $p^{+}$ prompts. Intuitively fine-tuning should learn to produce natural language thought from $p^{+}$. We do LORA with rank $8$, alpha $16$ and learning rate $2e^{-5}$ for 3 epochs. In figure 5 we compare finetuning results with activation steering, we showcase that finetuning InternLM2 yielded worse results than steering vectors. We attribute this behavior due to the limited number of $p^{+}$ prompts. We also made an attempt to fine-tune the model with the original set we sampled from LeanSTaR data.
Cost analysis We conduct all our experiments on Nvidia A100, for LORA finetuning, at rank 8 there are $1,338,656$ parameters. It took around 1 hour for training on our training data. In comparison steering only requires inference time updates making it much more efficient.
Cross Lingual Transfer To understand the true quality of these steering vectors we also evaluate InternLM2 on generating Coq proofs. This task helps us understand if the steering vectors only learn how to generate the natural language information while generating the proof or the core artifacts needed to imporve in Lean. In table 2 we show that steering vectors generate 1 valid proof in Coq from Putnambench. This cross lingual transfer validates the
Figure 5: Comparison between the pass rates achieved from models after doing LORA finetuning with parameters $\{rank=8,\alpha=16,learning\_rate=2\times 10^{-5}\}$ on prompts $p^{+}$. The pass rate are reported at 1x32x100 for InternLM2.
4.2.2 Layer responsibility
To localize where distilled informal thoughts most effectively influence proof generation, we perform layer-wise activation patching on our fine-tuned theorem-proving model. Specifically, we replace the model’s native activation vectors at contiguous layer blocks early (layers 5–13), middle (layers 14–24), and late (layers 25–30)—with those extracted under distilled reasoning prompts, and measure the resulting changes in miniF2F accuracy.
We find that interventions in the early layers yield only modest improvements, suggesting that these layers primarily encode low-level syntactic or symbol-manipulation features. Patching the middle layers leads to intermediate gains, indicating the emergence of conceptual inference patterns. Most notably, injecting distilled activation content into the final transformer blocks produces the largest accuracy boost. This highlights that the deepest layers provide the greatest leverage for steering proof-writing behavior, and that distilled reasoning signals most directly guide final proof synthesis and tactic sequencing.
        Selected Layers Pass Rates
        25-30 51.6%
        14-24 50.8%
        5-13 49.5%
Table 3: Pass rate at $2\times 32\times 600$ for InternLM2.5-StepProver at selected layers on miniF2f. This showcases how the proof construction relies on latter layers of language model and the intial layers are core part to comprehend and represent the theorems in activation space.
4.2.3 Avg Proof Length
To analyze the structural characteristics of the generated proofs, we track the usage of the have tactic, which is commonly used to introduce intermediate lemmas or helper statements. We find that proofs guided by informal reasoning contain a higher frequency of the have tactic, suggesting that such reasoning encourages a modular proof construction with explicitly stated intermediate steps.
Table 5 presents a comparative analysis of the theorems proved by InternLM2.5-StepProver with and without the use of a steering vector. The steering vector notably improves performance on shorter proofs, whereas the gains diminish for longer proofs. We hypothesize that this is due to the informal reasoning becoming less reliable as proof length increases as longer proofs require deeper logical consistency, and informal thoughts may introduce noise or misleading cues, thereby hindering accurate tactic prediction in such cases.
Model Decoding Random Steering
LLEMMA Sampling 22.7% 28.1%
Search 19.2% 26.3%
InternLM2 Sampling 21.4% 32.4%
Search 18.9% 26.8%
InternLM2.5-StepPtover Sampling 40.3% 52.4%
Table 4: Pass rates on miniF2F with randomized vectors and steering vectors.
Proof Length (steps) Without Steering With Steering
<5 76 83
6–10 16 13
10–15 10 11
15–30 15 13
30 < 11 8
Table 5: Proof length analysis on InternLM2.5-StepProver on miniF2F-test
4.2.4 Random Steering Vectors
A potential validity concern with any intervention involving activation patching is that the observed improvements might not stem from genuine performance enhancements but rather from activating backup circuits (McGrath et al.,, 2023; Lucchetti and Guha,, 2024). For instance, patching could merely introduce noise into the embedding space, inadvertently triggering alternative pathways that lead to the desired outcome. This phenomenon complicates the interpretability of both patches and steering vectors. To examine this, we conducted an experiment using a randomly generated steering vector (denoted ”Random” in Table 5). Our findings show that even random steering achieves a nonzero accuracy, albeit significantly lower than that of our computed steering vectors. We hypothesize that this residual accuracy arises due to backup circuits. Nevertheless, the substantially higher performance of our computed steering vectors suggests that our approach induces meaningful transformations toward the correct target.
4.3 How much does steering vectors improve proof writing?
We observed that steering helps to rerank responses which help us reduce the number of samples we expand at each node for Best First Search Tree method. In figure 6 we show how the pass rate improves with varying value of expansion width. This showcases the effectiveness of steering to improve
Figure 6: Pass at varying width (N) with best first search parameters $N\cdot 256\cdot 32$ in miniF2F-Test benchmark, we pick samples that are passed both with and without steering.

### v5

@@ 5 Results and Analysis
    Does steering enhance proof structure and search efficiency?
  4. 4.
    Are improvements consistent across different search budgets?
  5. 5.
    How effective and efficient are steering vectors compared to LoRA fine-tuning?
  6. 6.
    Do steering vectors generalize across provers?
5.1 Q1: Impact of Activation Steering
Activation steering consistently improves model performance across all tested configurations, enabling the discovery of proofs not derivable by the base models alone.
Model MiniF2F
Llemma 7B 26.6%
InternLM2-7B 28.7%
InternLM2.5-Step Prover 48.2%
LLEMMA-7B + Steering 28.1%
InternLM2-7B + Steering 32.4%
InternLM2.5-StepProver + Steering 66.4%
Table 1: Performance on MiniF2F (sampling decoding, 50×32×1).
Model PutnamBench
InternLM2 7B 4 (0.6%)
InternLM2.5-StepProver 6 (0.9%)
InternLM2 7B + Steering 4 (0.6%)
InternLM2.5-StepProver + Steering 7 (1.1%)
Table 2: Performance on PutnamBench (Lean). Results reported out of 657 attempts.
MiniF2F. As shown in Table 2, steering improves performance across all models on the miniF2F-Test benchmark (Zheng et al., 2022). We adopt the sampling decoding strategy from LeanStar (Lin et al., 2024), which mitigates variance from natural-language generation and provides more effective node selection than standard Best-First Tree Search (see Appendix for Best-First results).
Steering introduces structured natural-language comments that enhance mathematical reasoning and proof generation. Importantly, we find that steering enables a significant number of new proofs beyond the base model’s reach, suggesting that steering vectors guide the model toward otherwise inaccessible solution paths. Interestingly, models occasionally succeed despite incorrect informal reasoning, implying that steering exerts influence beyond surface-level natural language.
While steering increases successful proofs, it also introduces more failure cases. We hypothesize this stems from biases in the data used to construct steering vectors, which encode inductive priors that align well with some theorem classes but misalign with others. Despite this trade-off, the net effect is strongly positive, establishing activation steering as a lightweight, parameter-free alternative to fine-tuning.
PutnamBench. On PutnamBench (Table 2), steering improves success rates for both InternLM2 and InternLM2.5 on Lean, and achieves non-trivial gains on Rocq problems (details in Section A.3). These results highlight steering’s ability to support reasoning in more complex, multi-step proofs.
5.2 Q2: Layer Choice
We evaluate layer sensitivity following the strategy described in Section 3.3. Steering vectors are patched layer-wise to analyze their influence. Figure 4 and Table 4 show that later layers exert stronger influence on theorem-proving performance.
Selected Layers Pass Rate (%)
25–30 (Late) 51.6
14–24 (Middle) 50.8
5–13 (Early) 49.5
Figure 3: Pass rates at $2\times 32\times 600$ for InternLM2.5-StepProver on miniF2F.
Figure 4: Layer-wise ablation results.
These findings suggest that steering vectors primarily affect later-stage reasoning circuits. To ensure robustness, we also compare against random steering vectors, confirming that observed improvements arise from semantically meaningful directions rather than incidental noise.
5.3 Q3: Proof Structure and Search Efficiency
Table 3 shows that steering primarily benefits shorter proofs (<<5 steps), while gains diminish for longer ones (>>15 steps). Qualitative analysis suggests that informal reasoning introduces noisy intermediate steps for long proofs, reducing effectiveness. Interestingly, steering sometimes shortens proof length by guiding the model toward alternate proof strategies.
Proof Length Without Steering With Steering (total)
<<5 76 83 (94)
6–10 16 13 (19)
10–15 10 11 (20)
15–30 15 13 (18)
>>30 11 8 (11)
Table 3: Proof length analysis for InternLM2.5-StepProver on miniF2F.
While new proofs increase substantially, so do failure cases, likely reflecting inductive biases encoded in steering vectors. Characterizing these biases will be crucial for extending steering to broader proof distributions.
5.4 Q4: Search Budget
Figure 5 shows that steering yields consistent improvements across all search budgets $N$. Gains grow with larger budgets: at $N=100$, steering triples performance (6 vs. 2 proofs), and at $N=600$, steering delivers a 38% improvement (162 vs. 117 proofs). These results suggest that steering vectors not only boost baseline performance but also scale effectively with additional computational resources by guiding the search toward more productive proof trajectories.
Figure 5: Pass rates on miniF2F with varying search budgets.
5.5 Q5: Effectiveness vs. LoRA Fine-Tuning
We compare steering against LoRA fine-tuning (Hu et al., 2021) on InternLM2. LoRA models were trained on $p^{+}$ prompts with (rank=8, $\alpha_{\text{LoRA}}$=16) and (rank=32, $\alpha_{\text{LoRA}}$=64). Performance is reported on miniF2F after each epoch (Figure 6).
LoRA with higher rank outperforms steering, but lower-rank adaptations underperform, highlighting their limited representational capacity. In contrast, steering achieves competitive performance immediately, without additional training or parameters. This demonstrates that activation steering offers a highly parameter-efficient alternative, complementing but not replacing fine-tuning.
Figure 6: Comparison of steering and LoRA fine-tuning.
5.6 Q6: Generalization Across Provers
We evaluate if the steering vector is generalizable across different provers. In particular we use Putnambench (Tsoukalas et al., 2024) for evaluating cross lingual transfer for steering vectors. A particularly compelling finding emerges from our cross-system evaluation: despite the steering vectors being derived from Lean-based datasets, the augmented model successfully proves one Rocq problem an improvement over the base model’s zero success rate. Specifically, InternLM2.5 with steering successfully constructs a proof for Putnam 1988 B1 in Rocq, while the same model fails to complete the corresponding proof in Lean. In Appendix A.3 we showcase the proof written by InternLM2.5-StepProver.
This preliminary cross-system success suggests that steering vectors may capture reasoning patterns transferable across systems. The transferability suggests that our approach enables model to capture deeper mathematical concepts that transcend the specific formal language, encoding universal reasoning strategies that benefit theorem proving across different proof assistants. This finding has important implications for developing general-purpose mathematical reasoning systems, indicating that steering vectors trained on one formal system may provide value across the broader ecosystem of interactive theorem provers.